# Spatio-Temporal Context Prompting for Zero-Shot Action Detection

### v2

@@ 5.2 Zero-Shot Spatial-Temporal Action Detection
(a) Proposed components
Number mAP 80 11.77 100 12.85 150 12.46 200 11.46
(b) Interest tokens
@@
(c) Comparison with iCLIP
Rank mAP w/o LoRA 72.80 r = 2 64.52 r = 4 71.70 r = 8 74.55 r = 16 68.39
(d) LoRA in FFN
Person tokens mAP Without the adapter Patch-embedding 70.92 Image encoder 70.29 With the adapter 1-layer FC 67.37 2-layer FFN 74.55
(e) Person tokens
Temporal modeling mAP w/o temporal 68.03 average pooling 68.63 1-layer MHSA 74.55 2-layer MHSA 72.06
(f) Context tokens
Prompting mAP w/o prompting 66.98 only in last layer 66.02 in every layer 74.55
(g) Context prompting
Table 4: Ablation Study. We report frame mAP with 0.5 IoU threshold on the label split 1 of ZS-JHMDB and ZS-AVA.
5.3 Ablation Study
We present an ablation study in Table 4 to investigate different design choices in our method. The experiment in Table 4(b) is performed on the split 1 of ZS-AVA. Tabs. 4(c), 4(d), 4(e), 4(f) and 4(g) are conducted on the label split 1 of ZS-JHMDB, and Table 4(a) is performed on both.
Proposed components: We first investigate the importance of each component in Table 4(a). On J-HMDB, our proposed Person-Context Interaction effectively models the relationship between individuals and their surroundings, resulting in a 2.35 mAP improvement compared to the baseline. Furthermore, the Context Prompting module leverages context information to enhance text content, leading to an additional improvement of 7.57 mAP. On AVA, the above two components also demonstrate their effectiveness. Additionally, in multi-action videos, our Interest Token Spotting can identify context tokens most relevant to individual actions for prompting, further enhancing performance.
Interest tokens: In table 4(b), we conduct experiments using different numbers of interest tokens to observe their impact on the results. Our findings indicate that in multi-action videos, introducing too many tokens can potentially sample background noise unrelated to the action, thereby impacting the effectiveness of prompting.
Comparison with iCLIP [9]: We demonstrate the advantages of our method compared to [9] in Table 4(c). Without prompting, our method performs slightly better than [9], and we do not need to use additional object detectors and interaction blocks in the interaction modeling process. In addition, our prompting strategy uses multiple levels of context tokens to augment text content, resulting in an improvement of 7.57 mAP. However, the prompting method of [9] relies heavily on the results of interaction modeling, which limits their performance improvement.
LoRA in FFN: In Table 4(d), we investigate the impact of LoRA ranks. The results show that when we additionally train learnable matrices with rank 8, we can perform better than relying solely on CLIP’s pretrained weight.
Person tokens: Table 4(e) explores different ways of generating person tokens. Initially, the simple approach of pooling over the embeddings of all patches in the person crop fails to deliver satisfactory performance. Furthermore, equipping adapter can perform better than using person features at the image encoder level. This demonstrates that our adapter can effectively adapt person tokens, making them more suitable for the subsequent person-context interaction.
Context tokens: Table 4(f) shows the importance of temporal modeling. Using only the current frame to obtain context tokens results in the worst performance, indicating that aggregating temporal information is beneficial when identifying continuous actions. Additionally, equipping only a 1-layer MHSA can significantly improve performance compared to simple average pooling.
Context prompting: The results in Table 4(g) show the effectiveness of our prompting strategy. The results demonstrate that our prompting method, which utilizes tokens from low-level to high-level, yields better outcomes compared to using only high-level tokens from the last layer.
@@ 6 Conclusion
In this paper, we explore zero-shot spatio-temporal action detection. We propose a complete benchmark on J-HMDB, UCF101-24 and AVA. Besides, we propose a method to adapt the visual-language model for this task. The Person-Context Interaction employs pretrained knowledge to model the relationship between people and their surroundings, and the Context Prompting module utilizes the visual information to augment the text content. To address multi-action videos, we further introduce the Interest Token Spotting mechanism to identify the visual tokens most relevant to each individual action. The experiments demonstrate that our method achieves competitive performance compared to other video classification methods and can also handle multi-action videos.
@@
In the supplementary material, we present additional experimental results to substantiate the efficacy of our method and provide more details on the experimental settings. Initially, we elaborate our proposed benchmark for Zero-Shot Spatio-Temporal Action Detection in Sec 1. Subsequently, we showcase a visualization depicting interest tokens on ZS-AVA in Sec. 2. Then, in Sec. 3, we illustrate the distribution of text features on both ZS-JHMDB and ZS-UCF to assess the impact of prompting. We then provide the complexity analysis of our method and others in Sec. 4, and the results using ground-truth bounding boxes on each benchmark in Sec. 5. Finally, we discuss some limitations of our approach in Sec. 6, and give the implementation details of other methods in Sec. 7.
@@ 5 Results with Groundtruth Bounding Boxes
In this section, we present the results using ground-truth bounding boxes of the test data on ZS-AVA, ZS-JHMDB, and ZS-UCF in Tabs. 7, 8 and 9 respectively, to analyze the performance of all methods without localization errors.
@@
Frame mAP@0.5 Method split 1 split 2 split 3 split 4 avg Without the assumption of single-action video baseline 70.17 79.91 95.51 79.26 81.21 iCLIP [9] 71.06 76.57 93.44 74.98 79.01 ST-CLIP (Ours) 79.02 82.99 94.83 85.28 85.53 With the assumption of single-action video ActionCLIP [27] 74.92 83.33 89.46 80.58 82.07 A5 [12] 54.81 74.36 78.20 62.19 67.39 X-CLIP [18] 77.85 80.21 92.71 81.68 83.11 Vita-CLIP [28] 73.13 89.83 97.05 84.07 86.02 ST-CLIP (Ours) 85.02 87.87 97.33 90.27 90.12
@@
Frame mAP@0.5 Method split 1 split 2 split 3 split 4 avg Without the assumption of single-action video baseline 85.70 95.11 88.17 92.99 90.49 iCLIP [9] 91.30 89.91 81.56 79.12 85.47 ST-CLIP (Ours) 87.11 96.73 91.11 92.23 91.80 With the assumption of single-action video ActionCLIP [27] 91.33 94.39 85.26 87.90 89.72 A5 [12] 84.74 79.67 90.00 87.98 85.60 X-CLIP [18] 89.85 98.28 93.69 94.64 94.12 Vita-CLIP [28] 94.31 98.80 96.56 96.59 96.57 ST-CLIP (Ours) 92.90 98.75 93.97 94.44 95.02
@@
Tabs. 8 and 9 present the results on ZS-JHMDB and ZS-UCF. On ZS-JHMDB, our method has the best average performance whether using detected boxes or ground-truth boxes. On ZS-UCF, when we use ground-truth boxes to avoid false positives affecting the soft voting results, our method can achieve the second-best performance on most splits and on average. It is worth noting that the image quality of the UCF101-24 dataset is relatively blurry, with most of the characters occupying only a small part of the picture. This affects the accuracy of our individual classification, making it slightly disadvantageous compared to [28], which relies on the entire sampled frame.
@@ 7 Implementation Details
In this section, we provide the implementation details for each compared method in our experiments. For iCLIP [9], we employ the best-performing model featuring four types of interaction modules and Interaction-Aware Prompting. For ActionCLIP [27], to follow their approach, we sample 8 frames from each video, then utilize 6 transformer encoder layers to perform temporal modeling. Besides, we adopt the handcrafted prompts they proposed to prompt labels. For Efficient-Prompting [12], we sample 16 frames from each video, and employ the A5 model they proposed, which uses 2 encoder layers for temporal modeling and prepends/appends 16 vectors with the textual embeddings. For X-CLIP [18], we sample 8 frames from each video, and use the proposed Cross-frame Communication and 1-layer Multi-frame Integration Transformer to generate video features. Besides, we also leverage the Video-specific Prompting in their method to prompt labels. For Vita-CLIP [28], we sample 8 frames from each video, and utilize 8 Global Video-Level Prompts following their method. As for the training iterations, for iCLIP, we train the network for 7K iterations on ZS-JHMDB and 10K iterations on ZS-UCF as in their method. We train other video classification methods for 1K iterations on ZS-JHMDB and ZS-UCF, and more iterations will lead to lower performance due to overfitting. For all the aforementioned methods, we utilize the CLIP backbone and freeze the image and text encoder, consistent with our approach. We provide the details of training hyperparameters for all methods on ZS-AVA, ZS-JHMDB, and ZS-UCF in Tabs. 10, 11 and 12.
@@
Method Iterations Learning rate Warmup iterations Warmup factor Optimizer Batch size ActionCLIP [27] A5 [12] X-CLIP [18] 1000 0.0002 700 0.25 SGD 8 Vita-CLIP [28] iCLIP [9] 7000 0.0002 700 ST-CLIP (Ours) 3000 0.00025 800 0.25 SGD 8
@@
Method Iterations Learning rate Warmup iterations Warmup factor Optimizer Batch size ActionCLIP [27] A5 [12] X-CLIP [18] 1000 0.0002 1000 0.25 SGD 8 Vita-CLIP [28] iCLIP [9] 10000 0.0002 1000 ST-CLIP (Ours) 3000 0.00025 800 0.25 SGD 8

### v3

@@ 5.4 Ablation Study
Table 4: Proposed components
Comparison with iCLIP [10]: We demonstrate the advantages of our method compared to [10] in Table 6. Without prompting, our method performs slightly better than [10], and we do not need to use additional object detectors and interaction blocks in the interaction modeling process. In addition, our prompting strategy uses multiple levels of context tokens to augment text content, resulting in an improvement of 7.57 mAP. However, the prompting method of [10] relies heavily on the results of interaction modeling, which limits their performance improvement.
Context prompting: The results in Table 6 show the effectiveness of our prompting strategy. The results demonstrate that our prompting method, which utilizes tokens from low-level to high-level, yields better outcomes compared to using only high-level tokens from the last layer.
Method + prompt
@@
[…]
Table 5: Comparison with iCLIP
Prompting mAP
w/o prompting 66.98
only in last layer 66.02
in every layer 74.55
Table 6: Context prompting
Person tokens: Table 8 explores different ways of generating person tokens. Initially, the simple approach of pooling over the embeddings of all patches in the person crop fails to deliver satisfactory performance. Furthermore, equipping adapters can perform better than using person features at the image encoder level. This demonstrates that our adapter can effectively adapt person tokens, making them more suitable for the subsequent person-context interaction.
Context tokens: Table 8 shows the importance of temporal modeling. Using only the current frame to obtain context tokens results in the worst performance, indicating that aggregating temporal information is beneficial when identifying continuous actions. Additionally, equipping only a 1-layer MHSA can significantly improve performance compared to simple average pooling.
Person tokens mAP
Without the adapter
Patch-embedding 70.92
Image encoder 70.29
With the adapter
1-layer FC 67.37
2-layer FFN 74.55
Table 7: Person tokens
Temporal modeling mAP
w/o temporal 68.03
average pooling 68.63
1-layer MHSA 74.55
2-layer MHSA 72.06
Table 8: Context tokens
@@ 6 Conclusion
In this paper, we explore zero-shot spatio-temporal action detection. We propose a complete benchmark on J-HMDB, UCF101-24, and AVA. Besides, we propose a method to adapt the visual-language model for this task. The Person-Context Interaction employs pretrained knowledge to model the relationship between people and their surroundings, and the Context Prompting module utilizes visual information to augment the text content. To address multi-action videos, we further introduce the Interest Token Spotting mechanism to identify the visual tokens most relevant to each individual action. The experiments demonstrate that our method achieves competitive performance compared to other video classification methods and can also handle multi-action videos.
Acknowledgement: This work was supported by NVIDIA Taiwan Research & Development Center (TRDC).
@@
In the supplementary material, we present additional experimental results to substantiate the efficacy of our method and provide more details on the experimental settings. Initially, we elaborate our proposed benchmark for Zero-Shot Spatio-Temporal Action Detection in Sec. 1. Subsequently, we showcase a visualization depicting interest tokens on ZS-AVA in Sec. 2. Then, in Sec. 3, we illustrate the distribution of text features on both ZS-JHMDB and ZS-UCF to assess the impact of prompting. We then provide the complexity analysis of our method and others in Sec. 4. Besides, we provide more experimental analysis in Sec. 5 and describe how we handle multi-label prediction in Sec. 6. We also present the results using ground-truth bounding boxes on each benchmark in Sec. 7. Finally, we discuss some limitations of our approach in Sec. 8, and give the implementation details of other methods in Sec. 9.
@@ 4 Complexity Analysis
5 Additional Experimental Analysis
5.1 Interest Token Spotting on Single-Action Video
We further conduct experiments applying Interest Token Spotting on single-action video. We present the results on ZS-JHMDB label split 1 with detected boxes and ZS-UCF label split 1 with ground-truth boxes in Table 11. The results show that when a video contains only single action, all context tokens can be considered relevant to that action. Therefore, relying solely on the information from interest tokens during the prompting process is insufficient.
Prompting Tokens ZS-JHMDB ZS-UCF Interest Tokens 74.16 86.50 All Tokens 74.55 87.11
Table 11: Interest Token Spotting on single-action video
5.2 Complexity Analysis on Proposed Components
We present the complexity analysis on each of our proposed components in Table 12. The computational complexity (GFLOPs) is calculated on 1 AVA frame with 3 detected people, and we report the trainable parameters with ViT-L backbone. It is worth noting that the majority of Person-Context Interaction’s computational complexity arises from using CLIP’s pretrained image encoder to extract person features. Additionally, when applying Interest Token Spotting, we must prompt each person’s text features individually, rather than using a shared set of text features, which further increases the computational complexity.
Components GFLOPs Trainable params (M) Person-Context Interaction 583.04 22.97 Context Prompting 21.67 189 Interest Token Spotting 40.88 0
Table 12: Complexity Analysis on Proposed Components
5.3 Latency on Multi-Action Video
We report the latency of our ST-CLIP for processing multi-action videos in Table 13. We inference on 1 AVA video which contains 886 frames. Our method takes about 3 minutes for inference when using a single GPU.
Device Time (s) FPS 1 GPU 190.94 4.64 8 GPUs 40.22 22.03
Table 13: Latency on 1 AVA video
5.4 Additional Ablation Study
We present more ablation studies on the label split 1 of ZS-JHMDB and ZS-AVA.
LoRA in FFN: In Table 15, we investigate the impact of LoRA ranks. The results show that when we additionally train learnable matrices with rank 8, we can perform better than relying solely on CLIP’s pretrained weight.
Interest tokens: In table 15, we conduct experiments using different numbers of interest tokens to observe their impact on the results. Our findings indicate that in multi-action videos, introducing too many tokens can potentially sample background noise unrelated to the action, thereby impacting the effectiveness of prompting.
Rank mAP
w/o LoRA 72.80
r = 2 64.52
r = 4 71.70
r = 8 74.55
r = 16 68.39
Table 14: LoRA in FFN
Number mAP
80 11.77
100 12.85
150 12.46
200 11.46
Table 15: Interest tokens
6 Multi-Label Prediction on AVA
In the AVA dataset, each person must perform a pose action, and they may also engage in two additional types of actions: object interaction and person interaction. To handle multi-label prediction, after calculating the cosine similarity using person-context relational tokens and text features, we apply softmax as the activation function for pose actions, and sigmoid for the other two types — object interaction and person interaction.
@@ 7 Results with Groundtruth Bounding Boxes
In this section, we present the results using ground-truth bounding boxes of the test data on ZS-AVA, ZS-JHMDB, and ZS-UCF in Tabs. 16, 17 and 18 respectively, to analyze the performance of all methods without localization errors.
@@
Frame mAP@0.5 Method split 1 split 2 split 3 split 4 avg Without the assumption of single-action video baseline 70.17 79.91 95.51 79.26 81.21 ViCLIP [30] 54.30 81.32 63.44 73.55 68.15 iCLIP [10] 71.06 76.57 93.44 74.98 79.01 ST-CLIP (Ours) 79.02 82.99 94.83 85.28 85.53 With the assumption of single-action video ActionCLIP [29] 74.92 83.33 89.46 80.58 82.07 A5 [13] 54.81 74.36 78.20 62.19 67.39 X-CLIP [19] 77.85 80.21 92.71 81.68 83.11 Vita-CLIP [31] 73.13 89.83 97.05 84.07 86.02 ST-CLIP (Ours) 85.02 87.87 97.33 90.27 90.12
@@
Frame mAP@0.5 Method split 1 split 2 split 3 split 4 avg Without the assumption of single-action video baseline 85.70 95.11 88.17 92.99 90.49 ViCLIP [30] 75.21 74.79 44.07 58.22 63.07 iCLIP [10] 91.30 89.91 81.56 79.12 85.47 ST-CLIP (Ours) 87.11 96.73 91.11 92.23 91.80 With the assumption of single-action video ActionCLIP [29] 91.33 94.39 85.26 87.90 89.72 A5 [13] 84.74 79.67 90.00 87.98 85.60 X-CLIP [19] 89.85 98.28 93.69 94.64 94.12 Vita-CLIP [31] 94.31 98.80 96.56 96.59 96.57 ST-CLIP (Ours) 92.90 98.75 93.97 94.44 95.02
@@
Tabs. 17 and 18 present the results on ZS-JHMDB and ZS-UCF. On ZS-JHMDB, our method has the best average performance whether using detected boxes or ground-truth boxes. On ZS-UCF, when we use ground-truth boxes to avoid false positives affecting the soft voting results, our method can achieve the second-best performance on most splits and on average.
@@ 9 Implementation Details
In this section, we provide the implementation details for each compared method in our experiments. For iCLIP [10], we employ the best-performing model featuring four types of interaction modules and Interaction-Aware Prompting. For ActionCLIP [29], to follow their approach, we sample 8 frames from each video, then utilize 6 transformer encoder layers to perform temporal modeling. Besides, we adopt the handcrafted prompts they proposed to prompt labels. For Efficient-Prompting [13], we sample 16 frames from each video, and employ the A5 model they proposed, which uses 2 encoder layers for temporal modeling and prepends/appends 16 vectors with the textual embeddings. For X-CLIP [19], we sample 8 frames from each video, and use the proposed Cross-frame Communication and 1-layer Multi-frame Integration Transformer to generate video features. Besides, we also leverage the Video-specific Prompting in their method to prompt labels. For Vita-CLIP [31], we sample 8 frames from each video, and utilize 8 Global Video-Level Prompts following their method. For ViCLIP [30], we sample 8 frames from each video and use the video encoder to obtain the video feature map. Then, based on each person’s bounding box, we apply ROIAlign to extract individual person features for classification. As for the training iterations, for iCLIP, we train the network for 7K iterations on ZS-JHMDB and 10K iterations on ZS-UCF as in their method. We train other video classification methods for 1K iterations on ZS-JHMDB and ZS-UCF, and more iterations will lead to lower performance due to overfitting. For all the aforementioned methods except ViCLIP [30], we utilize the CLIP backbone and freeze the image and text encoder, consistent with our approach. As for ViCLIP [30], we freeze the text encoder and finetune the video encoder during training. The video encoder is pretrained on the InternVid-10M-FLT [30] dataset. We provide the details of training hyperparameters for all methods on ZS-AVA, ZS-JHMDB, and ZS-UCF in Tabs. 19, 20 and 21.
@@
Method Iterations Learning rate Warmup iterations Warmup factor Optimizer Batch size ActionCLIP [29] A5 [13] X-CLIP [19] 1000 0.0002 700 0.25 SGD 8 Vita-CLIP [31] ViCLIP [30] iCLIP [10] 7000 0.0002 700 ST-CLIP (Ours) 3000 0.00025 800 0.25 SGD 8
@@
Method Iterations Learning rate Warmup iterations Warmup factor Optimizer Batch size ActionCLIP [29] A5 [13] X-CLIP [19] 1000 0.0002 1000 0.25 SGD 8 Vita-CLIP [31] ViCLIP [30] iCLIP [10] 10000 0.0002 1000 ST-CLIP (Ours) 3000 0.00025 800 0.25 SGD 8